# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ 5.3. External Augmentation
External augmentation techniques have significantly boosted the effectiveness of LLMs across various security tasks, yielding competitive performance. We observed that only 28 out of the total 127 papers in LLM4Security, accounting for 22.05%, applied specific external augmentation techniques. From these studies, it is evident that external augmentation techniques have the potential to address issues such as hallucinations and high false positive rates caused by deficiencies in LLMs’ domain knowledge and task alignment. We believe that the integration of LLMs with external techniques will be a trend in the development of automated security task solutions.
@@
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 127 papers employing this approach. Fine-tuning techniques were used in 25.2% of the papers, while task-specific external augmentation techniques were adopted in 22.05% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks.
@@ 6.1. Data Collection
Data plays an indispensable and pivotal role in the training of LLMs, influencing the model’s capacity for generalization, effectiveness, and performance (Sun et al., 2022). Sufficient, high-quality, and diverse data areare imperative to facilitate the model’s comprehensive understanding of task characteristics and patterns, optimize parameters, and ensure the reliability of validation and testing. Initially, we explore the techniques employed for dataset acquisition. Through an examination of data collection methods, we classify data sources into four categories: open-source datasets, collected datasets, constructed datasets, and industrial datasets.
@@
Open-source datasets. Open-source datasets refer to datasets that are publicly accessible and distributed through open-source platforms or online repositories (Manocchio et al., 2024; Zhang et al., 2023f; Cai et al., 2024; Chen et al., 2023d). For example, the UNSW-NB15 dataset contains 175,341 network connection records, including summary information, network connection features, and traffic statistics. The network connections in the dataset are labeled as normal traffic or one of nine different types of attacks (Moustafa and Slay, 2015). The credibility of these datasets is ensured by their open-source nature, which also allows for community-driven updates. This makes them dependable resources for academic research.
@@
Collected datasets. Researchers gather collected datasets directly from various sources, such as major websites, forums, blogs, and social media platforms. These datasets may include comments from GitHub, harmful content from social media, or vulnerability information from CVE websites, tailored to specific research questions.
@@
Constructed datasets. The constructed dataset refers to a specialized dataset created by researchers through the modification or augmentation of existing datasets to better suit their specific research goals (Moskal et al., 2023; Wu et al., 2023a; Ahmad et al., 2023b; Koide et al., 2024). These changes could be made through manual or semi-automated processes, which might entail creating test sets tailored to specific domains, annotating datasets, or generating synthetic data. For instance, researchers might gather information on web vulnerabilities and the corresponding penetration testing methods, structure them into predefined templates to form vulnerability scenarios, and subsequently assess large language models using these scenarios (Deng et al., 2023a).
@@
Industrial datasets. Industrial datasets are data obtained from real-world commercial or industrial settings, typically consisting of industrial applications, user behavior logs, and other sensitive information (Zhang et al., 2023a; Liu et al., 2023b). These datasets are particularly valuable for research aimed at addressing real-world application scenarios.
@@
The Figure 6 illustrates the data collection strategies for LLM-related datasets. From the data depicted in the Figure 6, it can be observed that 47 studies utilize open-source datasets to train LLMs. The utilization of open-source datasets for training LLMs is predominantly attributed to their authenticity and credibility. These datasets are typically comprised of real-world data sourced from diverse origins, including previous related research, thereby ensuring a high degree of reliability and stability to real-world scenarios. This authenticity enables LLMs to learn from genuine examples, facilitating a deeper understanding of real-world security tasks and ultimately improving their performance. Additionally, due to the recent emergence of LLMs, there is indeed a challenge of the lack of suitable training sets. Hence, researchers often collect data from websites or social media and construct datasets to make the data more suitable for specific security tasks. We also analyzed the relationship between data collection strategies and the security domain. In certain domains such as network security, the preference for collecting datasets surpasses that of using open-source datasets. This indicates that obtaining data for applying LLMs to certain security tasks is still inconvenient. Among the 127 papers examined, only 2 studies utilized industrial datasets. This indicates a potential gap between the characteristics of datasets used in academic research and those in real-world industrial settings. This difference underscores the importance of future research exploring industrial datasets to ensure the applicability and robustness of large language models (LLMs) across academic and industrial domains. Some papers focus on exploring the use of existing LLMs, such as ChatGPT, in security tasks (Nair et al., 2023b, a). These papers often do not specify the datasets used for model training, as LLMs like ChatGPT typically do not require users to prepare their own training data for application scenarios.
@@ 6.2. Types of Datasets
The choice of data types plays a crucial role in shaping the architecture and selection of LLMs, as they directly influence the extraction of implicit features and subsequent decision-making by the model. This decision significantly impacts the overall performance and ability of LLMs to generalize (Liu et al., 2024). We conducted a thorough analysis and categorization of the data types utilized in LLM4Security research. Through examining the interplay between data types, model architectures, and task demands, our goal is to highlight the vital significance of data types in effectively applying LLMs to security-related tasks.
@@
Code-based datasets Vulnerable code 17 71 (Chen et al., 2023b) (Ferrag et al., 2023) (Chen et al., 2023c) (Chow et al., 2023)  (Liu et al., 2023d) (Thapa et al., 2022) (Zhang et al., 2023f) (Zhang et al., 2023g)  (Khare et al., 2023) (Ullah et al., 2023) (Liu et al., 2023c) (Wu et al., 2023a)  (Charalambous et al., 2023) (Alrashedy and Aljasser, 2024) (Ahmad et al., 2023a) (Chen et al., 2023d) (Ullah et al., 2024)
Source code 15 (Bakhshandeh et al., 2023) (Hu et al., 2023c) (Deng et al., 2023b) (Yang et al., 2023a)  (Xu et al., 2023) (Sun et al., 2023a) (Armengol-Estapé et al., 2024) (Pei et al., 2024)  (Song et al., 2022) (Liu et al., 2023c) (Sun et al., 2023b) (David et al., 2023)  (Hu et al., 2023a) (Gai et al., 2023) (Hu et al., 2024)
Bug-fix pairs 14 (Jin et al., 2023) (Li et al., 2023c) (Zhang et al., 2023d) (Yuan et al., 2022)  (Paul et al., 2023) (Olausson et al., 2024) (Xia and Zhang, 2022) (Huang et al., 2023)  (Zhang et al., 2023c) (Xia and Zhang, 2023b) (Xia and Zhang, 2023a) (Wang et al., 2023a)  (Zhang et al., 2023h) (Silva et al., 2024)
Bugs 7 (Li et al., 2023b) (Lee et al., 2023) (Sobania et al., 2023) (Paul et al., 2023)  (Huang et al., 2023) (Deng et al., 2023c) (Ahmad et al., 2023b)
@@
Code changes 3 (Yang et al., 2023c) (Du and Yu, 2023) (Wei et al., 2023c)
Vulnerability-fix pairs 2 (Zhang et al., 2023b) (Fu et al., 2022)
Bug fixing commits 2 (Zhang et al., 2023d) (Wang et al., 2023a)
@@
Subject protocol programs 1 (Meng et al., 2024)
@@
Text-based datasets Prompts 17 49 (Ali and Kostakos, 2023) (Moskal et al., 2023) (Deng et al., 2023a) (Temara, 2023)  (Charan et al., 2023) (Happe et al., 2023) (Fan et al., 2023) (Tol and Sunar, 2023)  (Zhang et al., 2023a) (Pearce et al., 2022) (Botacin, 2023) (Heiding et al., 2023)  (Roy et al., 2024) (Martin, 2023) (Scanlon et al., 2023) (Saha et al., 2023)  (Lin et al., 2023)
Log messages 6 (Liu et al., 2023b) (Chen et al., 2023e) (Karlsen et al., 2023) (Qi et al., 2023a)  (Han et al., 2023) (Shan et al., 2024)
Social media contents 5 (Hanley and Durumeric, 2023) (Hu et al., 2023b) (Cai et al., 2024) (Mets et al., 2023)  (Wang et al., 2023b)
Spam messages 4 (Labonne and Moran, 2023) (Moore et al., 2022) (Cambiaso and Caviglione, 2023) (Jamal and Wimmer, 2023)
Bug reports 3 (Fang et al., 2023) (Lee et al., 2023) (Du and Yu, 2023)
@@
Combined datasets Vulnerable code and vulnerability descriptions 2 2 (Liu et al., 2023d) (Chen et al., 2023c)
@@
Data type categorization. We categorize all datasets into three types: code-based, text-based, and hybrid data types. Table 9 provides a detailed breakdown of the specific data included in each category, derived from 127 studies. The analysis reveals that the majority of studies rely on code-based datasets, constituting a total of 71 datasets. This dominance underscores the superior code analysis capabilities of LLMs when trained for security tasks. These models demonstrate proficiency in understanding and processing code data, making them well-suited for security challenges such as vulnerability detection, program fuzzing, and traffic analysis. Their capacity to handle and learn from extensive code data enables LLMs to offer robust insights and solutions for various security applications.
Text datasets with numerous prompts (a total of 28) are commonly utilized for tasks lacking structured data, effectively guiding large language models (LLMs) through prompts to influence their behavior. While understanding the intricacies of training data might not be crucial for closed-source LLMs like ChatGPT, insights into data handling techniques for other models are still valuable. This is because black-box models can be fine-tuned with small-sized data inputs during usage. Among the 127 papers analyzed, text datasets rich in prompts are frequently used for training LLMs in security tasks, highlighting this trend. Additionally, specific security tasks necessitate particular text data inputs, such as system log analysis and harmful content detection.
@@
Data extraction Retrieve pertinent code segments from code-based datasets tailored to specific security tasks, accommodating various levels of granularity and specific task demands. Token-level, statement-level, class-level, traffic flow. (Sun et al., 2023a) (Chan et al., 2023)  (Montes et al., 2021)
Duplicated instance deletion Eliminate duplicate instances from the dataset to maintain data integrity and avoid repetition during the training phase. Removal of duplicate code, annotations, and obvious vulnerability indicators in function names. (Thapa et al., 2022) (Zhang et al., 2023f)  (Zhang et al., 2023h)
@@
Code representation Represented as tokens. Tokenize source or binary code as tokens. (Zhang et al., 2023g) (Xia and Zhang, 2022)  (Huang et al., 2023)
Data segmentation Divide the dataset into training, validation, and testing subsets for model training, parameter tuning, and performance evaluation. Partition the dataset based on specific criteria, which may include division into training, validation, or testing subsets. (Yang et al., 2023c) (Song et al., 2022)
The prevalence of vulnerable code (17), source code (15), and bug-fix pairs (14) in code-based datasets can be attributed to their ability to effectively meet task requirements. Vulnerable code naturally exhibits semantic features of code containing vulnerabilities to large language models (LLMs), thereby highlighting the distinguishing traits of vulnerable code when juxtaposed with normal code snippets. This aids LLMs in performing security tasks related to vulnerability detection. A similar rationale applies to bug-fix pairs. Source code serves as the backbone of any software project, encompassing the logic and instructions that define program behavior. Thus, having a substantial amount of source code data is essential for training LLMs to grasp the intricacies of programs, enabling them to proficiently generate, analyze, and comprehend code across various security tasks. Additionally, commonly used data types for bug fixes and traffic and intrusion detection, such as bugs (7) and traffic packets (4), are also widespread.
Some studies have utilized composite datasets containing multiple data types, such as vulnerable code and vulnerability descriptions. For instance, Liu et al. (Liu et al., 2023d) collected a dataset comprising CVE vulnerable code along with vulnerability descriptions and evaluated the performance of LLMs on vulnerability description mapping tasks based on this dataset.
6.3. Data Pre-processing
@@
Data preprocessing techniques for code-based datasets. We outline the preprocessing techniques utilized for code-based datasets, comprising five essential steps. Table 10 provides a comprehensive summary of each technique with examples. The initial step involves extracting data, retrieving relevant code snippets from diverse sources. Depending on the research task’s needs (Sun et al., 2023a; Montes et al., 2021), snippets may be extracted at different levels of detail, ranging from individual lines, methods, and functions to entire code files or projects. To prevent bias and redundancy during training, the next step removes duplicate instances by identifying and eliminating them from the dataset (Zhang et al., 2023f, h), enhancing diversity and uniqueness. Filtering follows, removing snippets that don’t meet predefined quality standards to ensure relevance to the security task and avoid noise (Ferrag et al., 2023; Yuan et al., 2022). Code representation converts snippets into suitable formats for LLM processing, often utilizing token-based representations for security tasks (Xia and Zhang, 2022). Finally, data splitting divides the preprocessed dataset into training, validation, and testing subsets (Yang et al., 2023c). Training sets train the LLM, validation sets tune hyperparameters, and testing sets assess model performance on unseen data. By adhering to these steps, researchers can construct structured code-based datasets, facilitating LLM application across various security tasks like vulnerability detection, program fuzzing, and intrusion detection.
@@
Data extraction Retrieve appropriate text from documentation based on various software engineering tasks. Attack description, bug reports, social media content, hardware documentation, etc. (Fayyazi and Yang, 2023) (Aghaei and Al-Shaer, 2023)  (Xia et al., 2024) (Hanley and Durumeric, 2023)  (Meng et al., 2023)
@@
Unqualified data deletion Delete invalid text data according to the specified rules. Remove certain symbols and words (rare words, stop words, etc.), or convert all content to lowercase. (Fang et al., 2023) (Cai et al., 2024)  (Aghaei et al., 2023b)
@@
Data segmentation Divide the dataset into training, validation, and testing subsets for model training, parameter tuning, and performance evaluation. Partition the dataset based on specific criteria, which may include division into training, validation, or testing subsets. (Rando et al., 2023) (Wang et al., 2023b)  (Liu et al., 2023e)
RQ4 - Summary (1) Based on different data sources, datasets are categorized into four types: open-source datasets, collected datasets, constructed datasets, and industrial datasets. The use of open-source datasets is the most common, accounting for approximately 38.52% in the 122 papers explicitly mentioning dataset sources. Collected datasets and constructed datasets are also popular, reflecting the lack of practical data in LLM4Security research. (2) We categorize all datasets into three types: code-based, text-based, and combined. Text-based and code-based types are the most commonly used types when applying LLMs to security tasks. This pattern indicates that LLMs excel in leveraging their natural language processing capabilities to handle text-based and code-based data in security tasks. (3) We summarize the data preprocessing process for different data types, outlining common data preprocessing steps such as data extraction, unqualified data deletion, data representation, and data segmentation.

### v4

@@ 5.3. External Augmentation
External augmentation techniques have significantly boosted the effectiveness of LLMs across various security tasks, yielding competitive performance. From these studies, it is evident that external augmentation techniques have the potential to address issues such as hallucinations and high false positive rates caused by deficiencies in LLMs’ domain knowledge and task alignment. We believe that the integration of LLMs with external techniques will be a trend in the development of automated security task solutions.
@@
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 185 papers employing this approach. Fine-tuning techniques were used in 31% of the papers, while task-specific external augmentation techniques were adopted in 36% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks.
@@ 6.1. Data Collection
Data plays an indispensable and pivotal role in the training of LLMs, influencing the model’s capacity for generalization, effectiveness, and performance (Sun et al., 2022). Sufficient, high-quality, and diverse data are imperative to facilitate the model’s comprehensive understanding of task characteristics and patterns, optimize parameters, and ensure the reliability of validation and testing. Initially, we explore the techniques employed for dataset acquisition. Through an examination of data collection methods, we classify data sources into four categories: open-source datasets, collected datasets, constructed datasets, and industrial datasets.
@@
Open-source datasets. Open-source datasets refer to datasets that are publicly accessible and distributed through open-source platforms or online repositories (Manocchio et al., 2024; Zhang et al., 2024b; Cai et al., 2024a; Chen et al., 2024b). For example, the UNSW-NB15 dataset contains 175,341 network connection records, including summary information, network connection features, and traffic statistics. The network connections in the dataset are labeled as normal traffic or one of nine different types of attacks (Moustafa and Slay, 2015). These are commonly used for benchmarking models or as a basis for fine-tuning on established tasks. The credibility of these datasets is ensured by their open-source nature, which also allows for community-driven updates. This makes them dependable resources for academic research.
@@
Collected datasets. Researchers gather collected datasets directly from various sources, such as major websites, forums, blogs, and social media platforms. These datasets may include comments from GitHub, harmful content from social media, or vulnerability information from CVE websites, tailored to specific research questions. Such data is often curated specifically for training or evaluating models on niche or newly defined security problems.
@@
Constructed datasets. The constructed dataset refers to a specialized dataset created by researchers through the modification or augmentation of existing datasets to better suit their specific research goals (Moskal et al., 2023; Wu et al., 2023a; Ahmad et al., 2024a; Koide et al., 2024). These changes could be made through manual or semi-automated processes, which might entail creating test sets tailored to specific domains, annotating datasets, or generating synthetic data. Generated synthetic data, for instance, often serves to augment training sets for fine-tuning LLMs when real-world data is scarce (as discussed further in Section 5). For instance, researchers might gather information on web vulnerabilities and the corresponding penetration testing methods, structure them into predefined templates to form vulnerability scenarios, which can then be used for evaluating LLM reasoning capabilities via prompting (Deng et al., 2024b).
@@
Industrial datasets. Industrial datasets are data obtained from real-world commercial or industrial settings, typically consisting of industrial applications, user behavior logs, and other sensitive information (Zhang et al., 2024e; Liu et al., 2023b). These datasets are particularly valuable for research aimed at training and evaluating models intended for real-world deployment and application scenarios.
@@
The Figure 6 illustrates the data collection strategies for LLM-related datasets. From the data depicted in the Figure 6, it can be observed that 73 studies utilize open-source datasets. The utilization of open-source datasets is predominantly attributed to their authenticity and credibility. These datasets are typically comprised of real-world data sourced from diverse origins, including previous related research, thereby ensuring a high degree of reliability and stability to real-world scenarios. This authenticity enables LLMs to learn from genuine examples, facilitating a deeper understanding of real-world security tasks and ultimately improving their performance when used for fine-tuning or evaluation. Additionally, due to the recent emergence of LLMs, there is indeed a challenge of the lack of suitable training sets. Hence, researchers often collect data from websites or social media (collected datasets) and construct datasets to serve as specific training, fine-tuning, or evaluation sets to make the data more suitable for specific security tasks. We also analyzed the relationship between data collection strategies and the security domain. In certain domains such as network security, the preference for collecting datasets surpasses that of using open-source datasets. This indicates that obtaining data for applying LLMs to certain security tasks is still inconvenient. Among the 185 papers examined, only 2 studies utilized industrial datasets. This indicates a potential gap between the characteristics of datasets used in academic research and those in real-world industrial settings. This difference underscores the importance of future research exploring industrial datasets to ensure the applicability and robustness of LLMs across academic and industrial domains. Some papers focus on exploring the use of existing LLMs, such as ChatGPT, in security tasks (Nair et al., 2023b, a). These papers often do not specify the datasets used for model training, as LLMs like ChatGPT typically do not require users to prepare their own training data for application scenarios but are instead guided via prompting or in-context learning.
@@ 6.2. Types of Datasets
The choice of data types plays a crucial role in shaping the architecture and selection of LLMs, as they directly influence the extraction of implicit features and subsequent decision-making by the model during both training and inference. This decision significantly impacts the overall performance and ability of LLMs to generalize (Liu et al., 2024a). We conducted a thorough analysis and categorization of the data types utilized in LLM4Security research. Through examining the interplay between data types, model architectures, and task demands, our goal is to highlight the vital significance of data types in effectively applying LLMs to security-related tasks.
@@
Code-based datasets Vulnerable code 24 90 (Chen et al., 2023b) (Ferrag et al., 2023) (Chen et al., 2023c) (Chow et al., 2023)  (Liu et al., 2023d) (Thapa et al., 2022) (Zhang et al., 2024b) (Zhang et al., 2023e)  (Khare et al., 2023) (Ullah et al., 2023) (Liu et al., 2023c) (Wu et al., 2023a)  (Charalambous et al., 2023) (Alrashedy and Aljasser, 2024) (Ahmad et al., 2023) (Chen et al., 2024b) (Ullah et al., 2024) (Cao et al., 2024) (Wen et al., 2024) (Chen and Jabbarvand, 2024) (Wu et al., 2024a)(Wu et al., 2024c) (Wang et al., 2024b) (Zhang and Zhang, 2024)
Source code 23 (Bakhshandeh et al., 2023) (Hu et al., 2023b) (Deng et al., 2023) (Yang et al., 2024a)  (Xu et al., 2023) (Sun et al., 2023) (Armengol-Estapé et al., 2024) (Pei et al., 2024)  (Song et al., 2022) (Liu et al., 2023c) (Sun et al., 2024) (David et al., 2023)  (Hu et al., 2023a) (Gai et al., 2023) (Hu et al., 2024a) (Fang et al., 2024a) (Lyu et al., 2024) (Eom et al., 2024) (Yu et al., 2024)(She et al., 2024) (Cai et al., 2024b) (Fang et al., 2024b) (Gohil et al., 2025)
Bug-fix pairs 16 (Jin et al., 2023) (Li et al., 2023c) (Zhang et al., 2023c) (Yuan et al., 2022)  (Paul et al., 2023) (Olausson et al., 2024) (Xia and Zhang, 2022) (Huang et al., 2023)  (Zhang et al., 2023b) (Xia and Zhang, 2024) (Xia and Zhang, 2023) (Wang et al., 2023b)  (Zhang et al., 2023f) (Silva et al., 2024) (Yin et al., 2024b) (Liu et al., 2024c)
Buggy code 10 (Li et al., 2023b) (Lee et al., 2023) (Sobania et al., 2023) (Paul et al., 2023)  (Huang et al., 2023) (Deng et al., 2024c) (Ahmad et al., 2024a) (Yang et al., 2024b) (Zhao et al., 2024) (Wang et al., 2024e)
@@
Code changes 3 (Yang et al., 2023b) (Du and Yu, 2023) (Wei et al., 2023c)
Vulnerability-fix pairs 3 (Zhang et al., 2023a) (Fu et al., 2022) (Zhang et al., 2024d)
@@
Subject protocol programs 1 (Meng et al., 2024a)
@@
Text-based datasets Prompts 19 60 (Ali and Kostakos, 2023) (Moskal et al., 2023) (Deng et al., 2024b) (Temara, 2023)  (Charan et al., 2023) (Happe et al., 2023) (Fan et al., 2023) (Tol and Sunar, 2024)  (Zhang et al., 2024e) (Pearce et al., 2022) (Botacin, 2023a) (Heiding et al., 2024a)  (Roy et al., 2024) (Martin, 2023) (Scanlon et al., 2023) (Saha et al., 2024)  (Lin et al., 2023) (Nazzal et al., 2024) (Liu et al., 2025)
Log messages 8 (Liu et al., 2023b) (Chen et al., 2023d) (Karlsen et al., 2023) (Qi et al., 2023a)  (Han et al., 2023) (Shan et al., 2024) (Jin et al., 2024) (Jeong et al., 2025)
Social media contents 7 (Hanley and Durumeric, 2023) (Hu et al., 2024b) (Cai et al., 2024a) (Mets et al., 2023)  (Wang et al., 2023c) (Thomas et al., 2024) (Lin et al., 2024)
Spam messages 4 (Labonne and Moran, 2023) (Moore et al., 2022) (Cambiaso and Caviglione, 2023) (Jamal et al., 2024)
Bug reports 4 (Fang et al., 2023) (Lee et al., 2023) (Du and Yu, 2023) (Feng and Chen, 2024)
Accident reports 2 (Guo et al., 2024) (Lu et al., 2024)
@@
Protocol Messages 2 (Meng et al., 2024b) (Wang et al., 2024d)
Program documentations 2 (Xia et al., 2024) (Wang et al., 2024f)
@@
Combined datasets Vulnerable code and vulnerability descriptions 7 23 (Liu et al., 2023d) (Chen et al., 2023c) (Nazzal et al., 2024) (Lin and Mohaisen, 2025) (Jiang et al., 2024) (Li et al., 2024b) (Chen et al., 2024a)
Buggy code and bug reports 4 (Zhang et al., 2024c) (Huang et al., 2024) (Wang et al., 2024c) (Tang et al., 2024)
Vulnerability-fix pairs and vulnerability descriptions 3 (Xu et al., 2024a) (Du et al., 2024a) (Ahmad et al., 2024a)
Bug-fix pairs and bug reports 3 (Chow et al., 2024) (Zhong et al., 2024) (Mündler et al., 2024)
Vulnerability-fix pairs and repair strategies 1 (Le et al., 2024)
Code changes and vulnerability descriptions 1 (Li et al., 2024b)
Source code and program documentations 1 (Zhang et al., 2024e)
Buggy code and commit messages 1 (Hossain et al., 2024)
Assertion code and hardware module descriptions 1 (Kande et al., 2024)
Decompiled code and variable names 1 (Xu et al., 2025)
@@
Data type categorization.  We categorize all datasets into three types: code-based, text-based, and combined data types. Table 8 provides a detailed breakdown of the specific data included in each category, derived from 185 studies. The analysis reveals that the majority of instances rely on code-based datasets, constituting a total of 90 instances. This dominance underscores the inherent code analysis capabilities of LLMs when applied for security tasks. These models demonstrate proficiency in understanding and processing code data, making code-based datasets suitable for fine-tuning models for tasks like vulnerability detection, program fuzzing, and bug repair. Their capacity to handle and learn from extensive code data enables LLMs to offer robust insights and solutions for various security applications either through fine-tuning or direct prompting.
Text-based datasets (total 60 instances) encompass a wide variety of textual information relevant to cybersecurity. While understanding the intricacies of training data might not be crucial for closed-source LLMs like ChatGPT, insights into data handling techniques for other models are still valuable. These datasets range from structured formats like prompts used for direct task execution (Deng et al., 2024b; Liu et al., 2023c) or logs used for fine-tuning anomaly detection models (Liu et al., 2023b; Chen et al., 2023d; Karlsen et al., 2023), to unstructured natural language found in bug reports (Fang et al., 2023; Du and Yu, 2023), CVE reports (Aghaei and Al-Shaer, 2023; Aghaei et al., 2023b), commit messages (Wu et al., 2024b; Zhang et al., 2024c), security policies (Nazzal et al., 2024), hardware documentation (Meng et al., 2023), and social media content (Hu et al., 2024b; Hanley and Durumeric, 2023). Specific security tasks necessitate particular text data inputs, often utilized for fine-tuning specialized classifiers (e.g., for harmful content) or providing context for LLM analysis via prompting (e.g., CVE reports).
The prevalence of vulnerable code (34 instances), source code (23), and bug-fix pairs (16) in code-based datasets can be attributed to their ability to effectively meet task requirements for fine-tuning or evaluation. Vulnerable code naturally exhibits semantic features of code containing vulnerabilities to LLMs, thereby highlighting the distinguishing traits of vulnerable code when juxtaposed with normal code snippets when training detection models. A similar rationale applies to bug-fix pairs, which are essential for fine-tuning program repair models. Source code serves as the backbone of any software project, encompassing the logic and instructions that define program behavior. Thus, having a substantial amount of source code data is essential for training LLMs (either during pre-training or fine-tuning) to grasp the intricacies of programs, enabling them to proficiently generate, analyze, and comprehend code across various security tasks. Additionally, commonly used data types such as buggy code (10) and traffic packages (4), are also widespread for training detection or analysis models.
Combined datasets, integrating multiple data types, represent a growing approach (23 instances identified). This strategy often pairs code artifacts with relevant natural language text, such as combining vulnerable code with vulnerability descriptions (Liu et al., 2023d; Chen et al., 2023c), buggy code with bug reports (Zhang et al., 2024c; Huang et al., 2024), or source code with program documentation (Zhang et al., 2024e). The key advantage lies in leveraging the LLM’s ability to process both structured code and unstructured text simultaneously (Wang et al., 2024a). By providing natural language context (e.g., descriptions of what a vulnerability entails or what a function should do), these combined datasets help LLMs achieve a deeper semantic understanding of the code during fine-tuning or prompting, leading to more accurate vulnerability detection, better code repair suggestions, or more effective code analysis compared to using code data alone.
6.3. Data Pre-processing
@@
Data extraction Retrieve pertinent code segments from code-based datasets tailored to specific security tasks, accommodating various levels of granularity and specific task demands. Token-level, statement-level, class-level, traffic flow. (Sun et al., 2023) (Chan et al., 2023)  (Montes et al., 2021)
Duplicated instance deletion Eliminate duplicate instances from the dataset to maintain data integrity and avoid repetition during the training phase. Removal of duplicate code, annotations, and obvious vulnerability indicators in function names. (Thapa et al., 2022) (Zhang et al., 2024b)  (Zhang et al., 2023f)
@@
Code representation Represent code as tokens, code graph, AST, etc. Tokenize source or binary code as tokens or transform code into call graph. (Zhang et al., 2023e) (Xia and Zhang, 2022)  (Huang et al., 2023) (Zhang and Zhang, 2024)
Data segmentation Divide the dataset into training, validation, and testing subsets for model training, parameter tuning, and performance evaluation. Partition the dataset based on specific criteria, which may include division into training, validation, or testing subsets. (Yang et al., 2023b) (Song et al., 2022)
@@
Data preprocessing techniques for code-based datasets. We outline the preprocessing techniques utilized for code-based datasets, often comprising five key steps. The initial step involves data extraction, retrieving relevant code snippets from diverse sources. Depending on the research task’s needs (Sun et al., 2023; Montes et al., 2021), snippets may be extracted at different levels of detail, ranging from individual lines, methods, or functions to entire code files or projects. To prevent bias and redundancy during training, the next step removes duplicated instances by identifying and eliminating them from the dataset (Zhang et al., 2024b, 2023f), enhancing diversity and uniqueness. Unqualified data deletion follows, removing snippets that don’t meet predefined quality standards (e.g., based on size, complexity, or presence of obvious hints) to ensure relevance to the security task and avoid noise (Ferrag et al., 2023; Yuan et al., 2022). Code representation converts snippets into suitable formats for LLM processing, typically involving tokenization for security tasks (Xia and Zhang, 2022). Finally, data splitting divides the preprocessed dataset into training, validation, and testing subsets (Yang et al., 2023b). Training sets are used to train the LLM, validation sets help tune hyperparameters, and testing sets assess model performance on unseen data.
@@
Data extraction Retrieve appropriate text from documentation based on various security tasks. Attack description, bug reports, social media content, hardware documentation, etc. (Fayyazi and Yang, 2023) (Aghaei and Al-Shaer, 2023)  (Xia et al., 2024) (Hanley and Durumeric, 2023)  (Meng et al., 2023)
@@
Unqualified data deletion Delete invalid text data according to the specified rules. Remove certain symbols and words (rare words, stop words, etc.), or convert all content to lowercase. (Fang et al., 2023) (Cai et al., 2024a)  (Aghaei et al., 2023b)
@@
Data segmentation Divide the dataset into training, validation, and testing subsets for model training, parameter tuning, and performance evaluation. Partition the dataset based on specific criteria, which may include division into training, validation, or testing subsets. (Rando et al., 2023) (Wang et al., 2023c)  (Liu et al., 2023e)
6.4. Data Augmentation using LLMs
A recurring challenge in applying LLMs to cybersecurity tasks is the scarcity of high-quality, labeled data. Datasets for vulnerabilities, malware, specific attack types, or other security events can be difficult to obtain in sufficient volume and diversity, hindering the training of robust models. To address this bottleneck, a number of studies have explored leveraging the generative capabilities of LLMs themselves for data augmentation (Dodge et al., 2020). The overarching goal is to enhance dataset size, diversity, and balance, thereby improving the robustness and generalization performance of security models trained (typically via fine-tuning) on the augmented data.
These LLM-based data augmentation strategies manifest in various forms depending on the task and data type. One common approach involves generating synthetic code samples used to enlarge fine-tuning datasets. This includes creating new examples of vulnerable code, such as RealVul using code transformations based on PHP vulnerability patterns (Cao et al., 2024), prompting LLMs with CWE details to generate pairs of vulnerable and non-vulnerable functions (Wang et al., 2024e), implanting specific bug types into correct code snippets to create debugging benchmarks (Zhong et al., 2024), or generating formally verified vulnerable and benign code snippets (Tihanyi et al., 2023). LLMs are also used to create specific malicious entities, for instance, generating diverse synthetic variations of malicious NPM packages based on known seeds to improve the fine-tuning of malware detection training (Yu et al., 2024). In the textual domain, LLMs synthesize relevant data often used for fine-tuning or as few-shot examples, such as examples of harmful online discourse (Martin, 2023), natural language repair suggestions (”oracles”) generated by prompting an LLM with vulnerable code and patches which then form a corpus for fine-tuning suggestion models (Zhang et al., 2024d), concise vulnerability explanations derived from patch features used for training auxiliary tasks during multi-task fine-tuning (Du et al., 2024a), or richer textual features like the explanatory ’debate text’ produced by EXPLAINHM to fine-tune a classifier for harmful meme classification (Lin et al., 2024). In a related application, LLMs assist in augmenting or constructing the knowledge bases essential for security tools, exemplified by VULTURE, which uses LLMs to analyze CVEs and commits for building a comprehensive third-party library vulnerability database (Xu et al., 2024a) used by the downstream detection tool.
RQ4 - Summary (1) Based on different data sources, datasets are categorized into four types: open-source datasets, collected datasets, constructed datasets, and industrial datasets. The use of open-source datasets is the most common, accounting for approximately 42% in the 173 papers explicitly mentioning dataset sources. Collected datasets and constructed datasets are also popular, reflecting the lack of practical data in LLM4Security research. (2) We categorize all datasets into three types: code-based, text-based, and combined. Code-based and text-based types are the most commonly utilized, with combined datasets representing a growing approach to integrate multiple modalities for enhanced semantic understanding. (3) We summarize the data preprocessing process applied to these various data types, outlining common steps such as data extraction, unqualified data deletion, suitable data representation, and data segmentation into training, validation, and testing sets. (4) To address the challenge of data scarcity in cybersecurity, a notable trend involves leveraging LLMs for data augmentation. Strategies include generating synthetic code samples (vulnerable or benign), specific malicious entities, diverse textual data (like security suggestions or explanations), and assisting in knowledge base construction.